Determine the profit or loss to call buyer and call writer for the following call options when the stock is selling at $32 just prior to expiration of the options and the option premium is $2.50.

a. $25 strike price

b. $30 strike price

c. $35 strike price
Answer #1

Option Premium = $2.50

Stock Price at Expiration = $32

a.

Strike Price = $25

Profit to Call Buyer = (32 - 25) - 2.50

Profit to Call Buyer = $4.50

Loss to Call writer = (25 - 32) + 2.50

Loss to Call Writer = -$4.50

b.

Strike Price = $30

Profit to Call Buyer = (32 - 30) - 2.50

Profit to Call Buyer = -$0.50

Loss to Call writer = (30 - 32) + 2.50

Loss to Call Writer = $0.50

c.

Strike Price = $35

Profit to Call Buyer = (0) - 2.50 (Option expires OTM)

Profit to Call Buyer = -$2.50

Loss to Call writer = (0) + 2.50 (Option expires OTM)

Loss to Call Writer = $2.50